40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 101 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  12459   Thu Sep 1 08:30:24 2016 SteveUpdateSUSITMY_UL is sick

So if the SRM satellite box is good, than the ITMY sensor UL or vacuum cabeling from sersor to sat amp is bad.

Quote:

Koji tweaked the alignment sliders till we were able to get the Y arm locked to green in a 00 mode, GTRY ~ 0.5 which is the prevent number I have in my head. The green input pointing looks slightly off in yaw, as the spot on the ITM looks a little misaligned - I will fix this tomorrow. But it is encouraging that we can lock to the green, suggests we are not crazily off in alignment.

[Ed by KA: slider values: ETMY (P, Y) = (-3.5459,  0.7050), ITMY (P, Y) =  (0.3013, -0.2127)]

While we were locked to the green, ITMY UL coil acted up quite a bit - with a large number of clearly visible excursions. Since the damping was on, this translated to somewhat violent jerking of ITMY (though the green impressively remained locked). We need to fix this. In the interest of diagnosis, I have switched in the SRM satellite box for the ITM one, for overnight observation. It would be good to narrow this down to the electronics. Since SRM is EQ-stopped, I did not plug in any satellite box for SRM. The problem is a difficult one to diagnose, as we can't be sure if the problem is with the LED current driver stage or the PD amplifier stage (or for that matter, the LED/PD themselves), and because the glitches are so intermittent. I will see if any further information can be gleaned in this regard before embarking on some extreme measure like switching out all the 1125 OpAmps or something...

Does anyone know if we have a spare satellite box handy? 

Is the spare sat amp is bad ?

Attachment 1: ITMY_SENSOR_UL.png
ITMY_SENSOR_UL.png
Attachment 2: spareSatAmp.jpg
spareSatAmp.jpg
  12458   Wed Aug 31 23:01:02 2016 gautamUpdateSUSthe chamber HEPA tents are back

Some more numbers we found while working in/around the chamber today:

Particle counts
Area 0.3 micron count (per cfm) 0.5 micron count (per cfm)
Vacuum chamber (vented) ITMY  ~5000 ~600
HEPA enclosure on ITMY door <100 0
Lab ~100,000 ~6000

These numbers were measured using our particle counter, which has a pump rate of 0.1 cfm, so the numbers above are 10x the numbers shown on the instrument after a measurement to account for this.

Essentially, the chamber is pretty dirty. Peeling the F.C with hard to reach optics like the ITM installed in place is not really feasible, and after peeling the F.C, we are looking at a best case of an additional 1-2 weeks in air to align the IFO, during which the optic is apparently exposed to quite a lot of particulates. In fact, with the high intensity flashlight left on, I actually saw some flecks of dust occassionally floating around inside the chamber while I was working on the optic. But this is just something we have to accept I guess.

  12457   Wed Aug 31 22:09:18 2016 gautamUpdateSUSY arm locked to Green

Koji tweaked the alignment sliders till we were able to get the Y arm locked to green in a 00 mode, GTRY ~ 0.5 which is the prevent number I have in my head. The green input pointing looks slightly off in yaw, as the spot on the ITM looks a little misaligned - I will fix this tomorrow. But it is encouraging that we can lock to the green, suggests we are not crazily off in alignment.

[Ed by KA: slider values: ETMY (P, Y) = (-3.5459,  0.7050), ITMY (P, Y) =  (0.3013, -0.2127)]

While we were locked to the green, ITMY UL coil acted up quite a bit - with a large number of clearly visible excursions. Since the damping was on, this translated to somewhat violent jerking of ITMY (though the green impressively remained locked). We need to fix this. In the interest of diagnosis, I have switched in the SRM satellite box for the ITM one, for overnight observation. It would be good to narrow this down to the electronics. Since SRM is EQ-stopped, I did not plug in any satellite box for SRM. The problem is a difficult one to diagnose, as we can't be sure if the problem is with the LED current driver stage or the PD amplifier stage (or for that matter, the LED/PD themselves), and because the glitches are so intermittent. I will see if any further information can be gleaned in this regard before embarking on some extreme measure like switching out all the 1125 OpAmps or something...

Does anyone know if we have a spare satellite box handy? 

  12456   Wed Aug 31 18:07:43 2016 JohannesUpdateSUSITMY free swinging

[Lydia, Gautam, Koji, Johannes]

Summary of things done today:

  • Rebalanced ITMY table
    • After waiting until today to see if the table would relax into a level position, engaged the earthquake stops for SRM and moved the large counterweight by ~4 inches. The table is now level to within ~0.1 mrad in direction of the access port
    • Since the relaxing seems to take some time, we will open ITMX and ETMX chamber tomorrow and level the tables with additional weights, so the springs can get used to 'levelness' again
  • Cleaned ITMY, SRM and SR2 optics
    • Koji drag-wiped all three optics and cleaned the table in general where accessible. He was able to remove the sliver discussed in elog https://nodus.ligo.caltech.edu:8081/40m/12455
    • We measured the particle count in the chamber and found it to be 4000 for 0.3 microns and 660 for 0.5 microns.
  • We pulled out stops on ETMY ITMY and roughly centered the OSEMs half-way, using photos of the previous OSEM rotation as a reference point for their orientation. We foudn that the green beam is hitting ITMY almost centered and that the reflection doesn't seem to steer off too much, but were not yet able to see any returned light on the ETMY cameras.

Unless we get lucky and get the green light to flash in the cavity by playing with the mirror alignment, we will open the ETMY chamber tomorrow. On one hand we can look for the reflected green light in the chamber, or alternatively the IR beam transmitted by ITMY. This way we can obtain estimates for the OSEM biasing and perform the final centering of the OSEMs. We will then also address the bounce mode minimization in ITMY and check if the previous orientations still hold.

  12455   Tue Aug 30 20:26:36 2016 gautamUpdateSUSITMY installed

[gautam, johannes, lydia]

Today we installed ITMY into position in the chamber.

  • First, we took the F.C coat off both faces
  • A stream of ionized nitrogen was used during the peeling process. We took as much care as possible not to blow towards the SRM. 
  • F.C. films came off smoothly. But when we looked at a picture we took prior to putting the optic in place, it looks like there may be a sliver of F.C. left on the optic. There are also a few specks of dust visible on the HR face, but well away from the clear aperture (see Attachment #1). Do we want to use isopropanol + optical tissue to try and remove these?
  • After F.C removal, we moved the optic into place against its stops. Returned OSEM connector tower to approximately its original place as it was moved to facilitate shifting the ITM to the edge of the table. 
  • I cleaned up the tangled mess of OSEM connector wires. On the ITMY tower, the OSEM cables have been tied using pieces of thin copper wire so as to avoid the wires straying into the beam path. Checked that wires are in grooves on both sides.
  • Unfortunately we were not able to start on setting up a cavity today, because when we checked the leveling of the ITM, we found that it was significantly not level. This is probably because the ITM was at the edge of the table. The cage is rather heavy and the location it was put in had a large lever arm. In any case, the table is slowly relaxing back to their usual state, Steve recommended we leave it overnight.
  • Other issues:
    • the UL sensor on ITMY also seemed to show some evidence of glitchy behaviour. Looking in the Satellite box, I didn't see any obvious probelms like I did for the ETMY box (for which I am not even sure if I did a legitimate fix anyways). I guess we have to keep observing and think about doing something about this if it really is problematic.
    • SRM barrel is pretty dusty. So is SR3. Do we want to clean these? If so how? F.C. or isoprop drag wipe?

We did some quick checks with the green beam and the IR beam. With the help of the custom Iris for the suspension towers, we gauged that both beams are pretty close to the center of the test mass. So we are in a not unreasonable place to start trying to align the beam. Of course we didn't check if the beam makes it to the ETM today.


The SRM OSEM sensor problem seems to have been resolved by moving the ITM back to its place as we suspected. The values are converging, but not to their pre-vent values (attachment #2). We can adjust these if necessary I guess... Or perhaps this fixes itself once the table returns to its neutral position. This remains to be monitored.


In the never-ending B-R mode reduction saga - we found what we think is an acceptable configuration now. Spectrum attached (Attachment #3). The top two OSEMs are now nearly 90 degrees rotated, while the bottom two are nearly horizontal. Anyways I guess we just have to trust the spectra. I should also point out that the spectra change rather significantly from measurement to measurement. But I think this is good enough to push ahead, unless anyone thinks otherwise?

Attachment 1: IMG_3052.JPG
IMG_3052.JPG
Attachment 2: SRM_sensor_level_comp.png
SRM_sensor_level_comp.png
Attachment 3: ETMY_BounceSpectra_30Aug2016.pdf
ETMY_BounceSpectra_30Aug2016.pdf
  12454   Tue Aug 30 15:44:28 2016 SteveUpdateSUS wire standoffs update

100 Sapphire prisms ordered. Delivery date 9-30-2016

~60 deg. prisms,

Size A=B=C=2 mm, length 5 mm

Material: sapphire

Surface quality: 5 micron RMS

Tolerances: +- 0.1 mm

This prism will be used as a mechanical component

No crystal orientation required

 

Quote:

There are some issues with 5 mm sapphire prism Atm5. It will cause  interference between one of the prisms and the Side OSEM.

Here are some drawings to see the issues with larger wire standoff.

The 2 mm prism will work.with a 1 mm longer dumbell.

Quotes requested from http://photomachining.com/laser-micromachining-photomachining-contact.html and http://www.optocity.com/ 

 

 

 

  12453   Tue Aug 30 09:51:37 2016 SteveUpdateSUSthe chamber HEPA tents are back

How to minimize particles entering the vacuum envelope.

 

Just the way it was in August 2011 vent and before.

The portable HEPAs were set up at ETMY and ITMY with CP STAT 100 curtains.

The  40m particles on the floor at ITMY 3000-5000 counts  of 0.5 micron cf / min and 0.3 micron size particles are 55,000 - 65,000 counts cf / min

At this condition the MET One Counter #3 on the floor inside the tent goes to zero count  of 0.5 micron and 20-40 counts cf / min  for 0.3 micron when the tent is slightly overpressured.

 

 

 

 

 

Attachment 1: ETMYtentHEPA.jpg
ETMYtentHEPA.jpg
Attachment 2: vented_for_60_days.png
vented_for_60_days.png
  12452   Mon Aug 29 21:53:14 2016 LydiaUpdateSUSETMY bounce mode coupling

[gautam, johannes, lydia]

We decided to try some different approaches on minimizing the ETMY bounce coupling today, since the peak height in the previously attched spectrum was  higher than the previously recorded levels in 2011 for all but the LR OSEM.

  • We recoreded a coarse reference spectrum with the light door off and the HEPA filter on. 
  • We attempted to match UL, UR and LL orientation to their apparent position in photos takes before the OSEMs were initially removed. We found that every OSEM showed a stronger bounce coupling, including those that had not been moved. We repeated the spectrum a few times, damping the optic and then turning the damping back off between each data set. The effect persisted, so we decided to move them back to where they started today (the final posiitons of our earlier optimization attempt). 
  • The peaks returned to approximately their reference levels for all except LL. UL was still the worst overall, so we attempted to more finely tune its angle without success: we saw increases in the peak height in both directions again (when turned approx. 3-4 degrees). 
  • We then turned the UL OSEM by 90 degrees. (See Attachment 1). Surprisingly, the 16.4 Hz peak height was greatly reduced (by a factor of 7-9). We took another spectrum to confirm and the peak was significantly lower than the reference yet again. We decided to leave the OSEM in this configuration in order to take a finer spectrum (See Attachment 2). However, this means that while the magnet looks well centered in the OSEM, the sideways range of motion is reduced and there are no earthquake stops preventing string side to side swings, so care should be taken in the ETMY chamber. 
  • Before starting, we had discussed a possilbe mechanism for bounce coupling that could potentially be minimized by turning the OSEM so that the nominal beam direction was horizontal. This would be possible if the beam of light inside the OSEM were directed toward the front or back, i.e. if it had some component parallel to the POS axis. Then, if the component of the beam in the plane of the optic's surface is oriented vertically, the bounce mode will move the edge of the magnet's shadow in and out of the center of the beam, allowing for strong coupling (proportional to the size of the angle). Turning the OSEM by 90 degrees would eliminate this kind of coupling. 
  • It's also true that (regardless of the mechanism) minimizing bounce couling for an OSEM maximizes the side coupling. The side coupling values of the pre-vent diagonalization matrices might be useful to compare to what we see now, to find out if the minimization we do in fact increases the side coupling (if not, maybe we are not truly minimizing the bounce coupling but observing some other effect, or our measurement method is too inconsistent.) Thoughts on this are welcome. 
  • Now that this improvement has been found, we should do a more thorough investigation to see where the true optimum position is (no small steps have been taken around this new position, and the minimum could also be at some angle in between).
  • In the interest of finishing the vent, should we try to find out exactly why the optimal angles seem so different than expected, or should we just try to minimize as best we can and move on to setting up the Y arm cavity? 
Attachment 1: IMG_3051.JPG
IMG_3051.JPG
Attachment 2: 42.png
42.png
  12451   Mon Aug 29 14:42:06 2016 gautamUpdateSUSETMY Satellite box diagnostics

I opened up the ETMY satellite box to investigate the glitches seen in the UL sensor output. 

Attachments #1 & 2: The connection to J4 from the satellite amplifier goes through a "satellite amplifier termination board", whose function, according to the schematic, is to prevent oscillations of the output amplifiers for the PD outputs. This seems to have been attached to the inside cover of the Satellite box by means of some sort of sponge/adhesive arrangement. The box itself gets rather hot however, and the sponge/adhesive was a gooey mess. I believe it is possible that some pins on the termination board were getting shorted - so if the 100 ohm resistor for the Ul channel that is meant to prevent the output amplifier oscillating was getting shorted, this could explain the problem.

For now, I cleaned off the old sponge/adhesive as best as I could, and used 4 pads of thick double sided tape (with measured resistance > 60Mohm) to affix the termination board to the inside of the box lid. In the ~3 hours since I have plugged the satellite box back in, there has been no evidence of any glitching. 

Of course, it could be that the problem has nothing to do with the termination board, and perhaps an OpAmp in the UL signal chain is damaged, but I stopped short of replacing these for now. I plan to push on with putting the IFO back together, and will keep an eye on this problem to see if more action is needed.

Also, if the inside of the ETMY satellite box had this problem of the sponge/adhesive giving way, it may be that something similar is going on in the other boxes as well. This remains to be investigated.

Attachment 1: IMG_6840.JPG
IMG_6840.JPG
Attachment 2: IMG_6841.JPG
IMG_6841.JPG
  12450   Mon Aug 29 09:37:05 2016 gautamUpdateSUSETMY UL sensor problematic

It looks like the problem is indeed in the Satellite box. Attachment #1 shows the second trend for the last 12 hours (~930pm 28 Aug 2016 - 930am 29 Aug 2016) for the ITMY and ETMY sensor signals. The satellite boxes for the two were switched during this time (the switch is seen at the leftmost edge of the plots). After the switch, ETMY UL has been well behaved, though ITMY UL shows evidence of excursions similar to what we have been seeing. All the ITMY coils are pulled out of the suspension cage currently, and are just sitting on the optical table, so they should just be reading out a constant value. I think this is conclusive evidence that the problem is with the Satellite box and not the OSEM itself. I will pull the Satellite box out and have a look at its innards to see if I can find the origin of the problem...

Attachment 1: satelliteBox.png
satelliteBox.png
  12449   Sun Aug 28 21:40:11 2016 gautamUpdateSUSETMY UL sensor problematic

I wanted to observe the UL coil for any excursions over the weekend. Looking at the 2 day trend, something is definitely wrong. These glitches/excursions are much more pronounced than what is seen in the pre-vent plots Steve had put up.

In order to try and narrow down whether the problem is with the Satellite box or the LED/PD themselves, I switched the Satellite box at the Y end with the Satellite box for ITMY (at ~930pm tonight). Hopefully over a 12 hour observation period, we see something that will allow us to make some conclusion. 

Attachment 1: ETMY_UL_problematic.png
ETMY_UL_problematic.png
  12448   Fri Aug 26 17:48:14 2016 gautamUpdateSUSbounce mode coupling reduction

We worked on reducing the bounce mode coupling into the sensor signals today. After some trial and error, essentially following the procedure I had put up in my previous elog, we think we were successful in reducing the coupling. We have now left the optic free swinging, so that we can collect some data and look at a spectrum with finer bandwidth. But as per the methodology we followed, we saw that the peak height corresponding to the bounce mode increased when we rotated the OSEM either side of its current position (except for the side OSEM, which we felt was in a good enough position to warrant not touching it and messing it up - of course only the spectrum will tell us if we are right or not. I also took some pictures with the camera with the IR filter removed, but we couldn't get any real information from these photos. I also checked with Jenne and Jamie who both suggested that they didn't have any metric with which they judged if the rotation of the OSEM was good enough or not. So we will wait to have a look at the spectrum from later tonight, and if it looks reasonable enough, I vote we move on. As Eric suggested, perhaps we can repalce the UL OSEM coil and see if that solves the apparent UL coil problem. Then we should move on to putting the arm cavity together.

Addendum 11pm 26 Aug 2016: I've uploaded the spectra - looks like our tweaking has gained us a factor of ~2 on LL, LR and SD, and no significant improvement on UL and UR compared to yesterdays spectrum.

Attachment 1: ETMY_BounceSpectra_26Aug2016_1.pdf
ETMY_BounceSpectra_26Aug2016_1.pdf
  12447   Fri Aug 26 14:09:46 2016 ericqUpdateSUSETMY UL sensor problematic

We do indeed have a box of clean spare OSEMs, it should be out with all of the other boxes of clean stuff we had for the suspension building. You could also try swapping in a different satellite box, to see if the circuit powering the OSEM PD is to blame.

  12446   Fri Aug 26 14:05:38 2016 SteveUpdateSUSETMY UL sensor problematic

This problem has existed well before the vent

 

Attachment 1: ETMY-ULhistory.png
ETMY-ULhistory.png
  12445   Fri Aug 26 11:35:44 2016 gautamUpdateSUSETMY UL sensor problematic

I've been noticing that the ETMY UL sensor output has been erratic  over the last few days. It seems to be jumping around a lot, even though there is no discernable change in any of the other sensor signals. Damping is OFF, which means the sensor signals should just be a reflection of actual test mass motion. But the fact that only one sensor output is erratic leads me to believe that the problem is in the electronics. I've also double checked that we aren't touching any EQ stops. Also, we had centered all the sensor outputs to half their maximum value pretty carefully. But looking at the Striptool traces, I now find that the UL sensor output has settled at some other value. Simply removing the OSEM connector and plugging it in again leads to the sensor output going back to the carefully centered value. Could it be that the photodiode has gone bad? If so, do we have spare OSEMs to use? I will also re-squish the satellite box cables to see if that fixes the problem.

Attachments:

Attachment #1: Sensor output spectra around the bounce mode peak. Nothing was touched inside the chamber between the time this spectrum was taken and the spectrum I put up last night (in fact the chamber was closed)

Attachment #2: UL sensor output is erratic, while the others show no glitching. This supports the hypothesis that the problem is electronic. The glitch itself happened while the chamber was closed.

Attachment #3: The only difference between this trace and Attachment #2 is that the UL connector was removed and plugged in (OSEM wasn't touched)

Attachment 1: ETMY_BounceSpectra_26Aug2016.pdf
ETMY_BounceSpectra_26Aug2016.pdf
Attachment 2: 41.png
41.png
Attachment 3: 19.png
19.png
  12444   Thu Aug 25 21:11:43 2016 gautamUpdateSUSETMY back in IFO

There was some confusion as to the order in which we should go about trying to recover the Y arm. But here are the steps we decided on in the end.

  1. Use the tip tilts to make sure the input beam is hitting roughly the center of ETMY, with ITMY left out.
  2. Use the reflected beam from the ETM as viewed in the ITM chamber to set the pitch bias on ETM.
  3. Center OSEM coils on ETM, rotate them to minimize bounce mode coupling into the sensor signals.
  4. Install the ITM, look for cavity flashes, and use alignment biases to try and lock the Y arm in air.

Yesterday, Eric, Johannes and I tried to do step 1, but after some hours of beam walking, we were unsuccessful. Today morning, Koji suggested that the ITM wedge could be playing a part - essentially, over 40m, the wedge would shift the beam horizontally by ~30cm, which is kind of what we were seeing yesterday. That is, with 0 biases to the tip tilts, we could find the beam in the ETM chamber, towards the end of the table, ~30cm away from where it should be (since the input pointing is adjusted taking this effect into account, but we were doing all of our alignment attempts without the ITM in).

So, we shifted strategy today. The idea was to trust that the green beam was well aligned to the cavity axis (we had maximized the green transmission before the vent), and set the pitch bias voltage to ETMY by making the reflected beam overlap with itself. This was done successfully, and we needed to apply a pitch bias of ~-2.70 (value on the MEDM screen slider), which agrees well with what I was seeing in the cleanroom. We then adjusted the OSEMs to bring the sensor outputs to half their nominal maximum value. Next, we went into the ITMX chamber, and were able to find the green beam, at the right height, and approximately where we expect the center of the ITM to be (this supports the hypothesis that the green input pointing was pretty good). I am however concerned if this is truly the right value of the bias for making a cavity with the ITM, because the pre-vent value of the pitch bias slider for ETMY was at -3.7, which is a 30% difference from the current value (and I can't think of a reason why this should have changed, the standoffs weren't touched for ETMY). If we go ahead and fine tune the OSEMs rotationally assuming this is the right bias to have, we may end up with sub-optimal bounce mode coupling into the sensor signals if we have to apply a significantly larger/smaller offset to realise a cavity? The alternative is to put in the ITM, and set the pitch balance using the IR beam, and then go about rotating OSEMs. The obvious downside is that we have to peel the F.C. off, risking dirtying the ITMs.

For much of the rest of the day, we were trying to play with the rotation of the OSEM coils in order to minimize the bounce mode coupling into the sensor signals. We weren't able to come up with a good scheme to do this measurement, and I couldn't find any elog which details how this was done in the past. The problem is we have no target as to how good is good enough, and it is extremely difficult to gauge whether our rotation has improved the situation or not. For instance, with no rotation of the OSEMs, by observing the bounce mode peak height over a period of 20-30 minutes, we saw the peak height change by a factor of at least 3. This is not really surprising I guess, because the impulses that are exciting the bounce mode are stochastic (or at least they should be), and so it is very hard to make an apples to apples comparison as to whether a rotation has improved the situation on.

After some thought, the best I can come up with is the following. If anyone has better ideas or if my idea is flawed, or if this is a huge waste of time, please correct me!

  1. Adopt this spectrum (except the side signal) as a reference for what constitutes "good" rotational orientation of the OSEMs (even though it is for ETMX not ETMY).
  2. Start with one coil. The suspension assembly document tells us to expect the orientation with minimal bounce coupling to be located within 20 degrees of "the vertical", the vertical being defined as that orientation in which the line connecting the LED and PD as seen by eye is vertical. So start with the coil oriented vertically, as best as possible by eye.
  3. Damp the optic for ~1min, with the curtain covering the chamber entrance. Ideally, we want the door back on, as this lowers the noise floor significantly, but it is too cumbersome to replace even the light door so I suppose we will have to compromise.
  4. Take a reference spectrum. In the interest of time, I think a bandwidth of 0.1Hz on the Fourier Transform should be sufficient. (Tangentially related - the BW you specify in the measurement setup in DTT doesn't seem to be the BW with which the spectrum is computed, I wonder why that is?)
  5. It is basically impossible to rotate the coil continuously. So divide the range to be explored into steps (so each step will involve rotating the coil by ~2 degrees (I don't know if this number is physically feasible, but some discrete step will be involved). Rotate the coil, center it such that the sensor output is close to half the maximum.
  6. Pull the curtain down, damp the optic, and take another spectrum. If the bounce mode peak is higher, abandon this direction of rotation, and rotate the other way. We accept as the optimal position the one from which the bounce mode peak height gets worse by rotating to either side.

Of course, this method assumes that the excitation into the bounce mode is a constant over time. I'm also attaching the spectrum of the OSEM sensor signals right now - the optic is in the chamber, free swinging (no damping) with the door on (so it is fairly quiet). The LR signal seems to be the best (indeed seems to match the levels in this plot), but it is not clear whether the others can be improved or not.


There was also some concern as to whether we will be able to see the beam in the ETMX chamber once the ITM has been re-installed. Assuming we get 100mW out of the IMC, PRM transmission of 5.5%, and ITM transmission of 1.4%, we get ~35uW incident on the ETM, which while isn't a lot, should be sufficient to see using an IR card.

Attachment 1: ETMY_BounceSpectra_25Aug2016.pdf
ETMY_BounceSpectra_25Aug2016.pdf
  12443   Thu Aug 25 20:07:35 2016 gautamUpdateSUSOSEM issues - maybe resolved?

[lydia, johannes, gautam]

While struggling to minimize the bounce mode coupling into the sensor signals, we briefly poked into the ITMY chamber, and think that we understand the origin of the problem, at least for the SRM.

Essentially, we believe that moving the ITM from its nominal position to the edge of the table has shifted the table leveling such that the optic (SRM) is tilted backwards (hence the magnets are completely occluding the LEDs) and that perhaps the optic is in contact with one or more of the bottom EQ stops (hence the signal is stationary, no oscillations visible. The timing of the signals going dark as Eric mentioned supports this hypothesis. The reason why we believe this to be the case is that when I was trying to loosen the screw on the clamp holding the ITMY cage to the table, we saw ~1Hz signals from all 5 SRM OSEM sensors, though they were well away from the nominal equilibrium values. The arrangement of towers in the chamber right now did not permit me to get a good look at the SRM magnets, but I believe they are all still attached to the optic, and that they are NOT stuck to the OSEM coils. If this is indeed the case, putting ITMY back in will solve the issue completely.

It is not clear what has happened to the LR coil on the PRM - could it be that during the venting process, somehow the LR magnet got stuck to the OSEM? If so, can we free it by the usual bias jiggling?

  12442   Thu Aug 25 19:03:56 2016 PrafulUpdateElectronicsAcoustic Tab and Amp Suspension

My box has been suspended in the PSL using surgical tubing, and it has been connected to C1:PEM-MIC_1 (C17) with a BNC. I made a braided power cable as well but it turned out to be slightly too short... Once this is fixed, everything should be ready and we can see if it's working correctly. I also set up a new tab on the summary pages for this channel:

https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1154941217-1154942117/pem/acoustic/

This data is back from when I had my solderless breadboard running near MC2. I'll add this tab to the real pages once the box is working (which could be a while since I'm gone for a month). Let me know if you see any issues with either the tab or the box/cables.

  12441   Thu Aug 25 15:43:29 2016 ericqUpdateSUSOSEM issues

We've seen for some time now that one of the PRM OSEM signals has been gone, and all of the SRM signals seem dark. We had tried squishing various cables to no avail.

Today I played some "musical satellite boxes," in an attempt to see if the problems are in the chambers or in the signal chains. That is, I swapped the OSEM cables from the vacuum feedthroughs between the satellite boxes, and observed what happened.

It seems clear that something is up with SRM inside the chamber. For PRM, it's not so clear...

  PRM Satellite SRM Satellite BS Satellite
PRM OSEMs LR low + small fluctuations LR low + small fluctuations LR low + small fluctuations
SRM OSEMs No signals No Signals No Signals
BS OSEMs LR low, flat OK OK

Somehow, issues with the LR channel follow both the PRM OSEMs and the PRM satellite box. frown

PRM LR first went dark on Jul 2nd, after the IFO was vented, but before we took any doors off (which happened on the 5th). I'm not sure what may have caused this.

SRM OSEMS first went dark on the evening of Jul 18, the day before ELOG 12310, when ITMY was moved in the same chamber. Maybe this ELOG was written about work the day before, but the sensors show disturbances over the course of hours. I think we need to double check the connections in chamber. 

  12440   Thu Aug 25 08:19:25 2016 Max IsiUpdateGeneralSummary pages down due to cluster maintenance

The system is back from maintenance and the pages for last couple of days will be filled retroactively by the end of the week.

Quote:

Summary pages down today due to schedulted LDAS cluster maintenance. The pages will be back automatically once the servers are back (by tomorrow).

 

  12439   Wed Aug 24 23:47:30 2016 PrafulUpdateElectronicsFinished Prototype Box

Gautam helped me drill holes in a metal box and I set up my circuit inside. Everything seems to be working so far. Tomorrow I'll be suspending the box near the PSL and setting up a data channel. Attached are some pictures of the box- sorry some of the angles turned out weird.

Attachment 1: out1.pdf
out1.pdf
Attachment 2: out2.pdf
out2.pdf
Attachment 3: out3.pdf
out3.pdf
Attachment 4: in1.pdf
in1.pdf
Attachment 5: in2.pdf
in2.pdf
  12438   Wed Aug 24 19:37:55 2016 KojiUpdateElectronicsDecoupling capacitor 101

Yes

Interesting articles how they should only be used for power decoupling and not in the signal path.

http://www.edn.com/design/analog/4416466/Signal-distortion-from-high-K-ceramic-capacitors

http://www.edn.com/design/analog/4426318/More-about-understanding-the-distortion-mechanism-of-high-K-MLCCs

  12437   Wed Aug 24 14:44:33 2016 PrafulUpdateElectronicsDecoupling capacitor 101

Do these look good for the ceramic capacitors? We're running low.

http://www.mouser.com/ProductDetail/Vishay-BC-Components/K104K15X7RF53L2/?qs=sGAEpiMZZMuMW9TJLBQkXmrXPxxCV7CRo6C15yUYAos%3d

Quote:

What I suggested was:
- For most cases, power decoupling capacitors for the regulators should be ~100nF "high-K ceramic capacitors" + 47uF~100uF "electrolytic capacitors".
- For opamps, 100nF high-K ceramic should be fine, but you should consult with datasheets.
- Usually, you don't need to use tantalum capacitors for this purpose unless specified.
- Don't use film capacitors for power decoupling.

79XXs are less stable compared to 78XXs, and tend to become unstable depending on the load capacitance.
One should consult with the datasheet of each chip in order to know the proper capacitors values.
But also, you may need to tweak the capacitor value when necessary. Above recipe works most of the case.

 

  12436   Wed Aug 24 14:11:09 2016 PrafulUpdateElectronicsMicrophone Testing

I added an EM172 to my soldered circuit and it seems to be working so far. I have taken a spectra using the EM172 in ambient noise in the control room as well as in white noise from Audacity. My computer's speakers are not very good so the white noise results aren't great but this was mainly to confirm that the microphone is actually working.

white_v_ambient.pdf

Attachment 1: white_v_ambient.png
white_v_ambient.png
Attachment 2: white_v_ambient.pdf
white_v_ambient.pdf
Attachment 3: white_v_ambient.pdf
white_v_ambient.pdf
  12435   Tue Aug 23 22:58:16 2016 KojiUpdateElectronicsDecoupling capacitor 101

What I suggested was:
- For most cases, power decoupling capacitors for the regulators should be ~100nF "high-K ceramic capacitors" + 47uF~100uF "electrolytic capacitors".
- For opamps, 100nF high-K ceramic should be fine, but you should consult with datasheets.
- Usually, you don't need to use tantalum capacitors for this purpose unless specified.
- Don't use film capacitors for power decoupling.

79XXs are less stable compared to 78XXs, and tend to become unstable depending on the load capacitance.
One should consult with the datasheet of each chip in order to know the proper capacitors values.
But also, you may need to tweak the capacitor value when necessary. Above recipe works most of the case.

  12434   Tue Aug 23 19:35:38 2016 gautamUpdateSUSETMY back in IFO

[johannes, gautam]

Summary: Today we moved the suspended ETMY optic back into the chamber from the cleanroom. Once in the chamber, we positioned the optic using the stops that marked the previous position of the optic. We then shortened the arm length by 19mm (in order to match the X and Y arm lengths. The F.C. coat on the HR face was removed prior to the final placement of the optic. We then adjusted the OSEM positions in their holders to get the sensor outputs to half their maximum value.

We did not get to check where the input beam hits the optic or see if the pitch balance of the optic is such that the reflected beam makes it back to the ITM. The plan for tomorrow is to do this. 


Part 1: Cleanroom work

  • We worked a little more on trying to adjust the rotational position of the OSEM coils in order to minimize the coupling of the bounce mode into the sensor signals. 
  • We had limited success in this regard. After about an hour, we concluded that it made more sense to do this in the chamber itself. For one thing, the drive electronics for the Y end are different (in the cleanroom, we are using the X end electronics, satellite box etc.).
  • We adjusted the position of the OSEMs till the sensor output readout was half the open value as best as we could. We also made sure that the wire was in the groove on both sides and that the magnets were well centered in the vertical direction relative to the OSEM coils and that there was no danger of knocking any magnets off (see attached pictures).
  • We then engaged all the EQ stops, and transferred the suspension cage to a cart (topped with Al foil, wiped clean) for transportation to the Y-end (with OSEMs left in).

Part 2: Transportation of optic

  • Nothing special here, just took great care while going over bumps near doors between the cleanroom and the IFO, and along the Y-arm itself.
  • Definitely a 2 man job - one person can lift a pair of wheels over any bumps while the other can make sure there is no danger of the cage toppling over. 

Part 3: Chamber work

  • PSL shutter was closed for this part of work. Earlier today, I found that C1SUSAUX had failed yet again (why are all the slow computers dying more often nowadays?!). I restarted the slow machine, and locked the mode cleaner. The alignment hadn't drifted so much from when EricQ had last aligned the IMC, and with only minimal tweaking, I was able to lock the IMC and see a beam on the REFL camera.
  • First, I transferred the suspension cage onto the edge of the table inside the chamber. Care was taken not to accidentally place the cage onto the trailing OSEM wires.
  • There were some specks of dust on the barrel of the optic, and also the cage. These were removed with clean wipes and isopropanol.
  • I judged that it would be too precarious to remove the F.C. with the optic in its final desired position. So we decided to take the coat off with the optic at the edge of the table. The central part of the HR face looks pretty clean. Even though the whole HR face was cleaned with F.C., the part that was left uncovered prior to putting the optic back into the chamber has a few specks of dust on it (see attachments). These could not be removed just by blowing ionized air. I was hesitant to drag wipe the optic, so I left things as is. In any case, the optic as a whole is MUCH cleaner (to my eye at least) than prior to the cleaning. 
  • Conveniently, the stops marking the previous position of the optic were on the far side and back.
  • Since we wanted to shorten the Y arm length by 2 cm, we placed a clean steel ruler of width 19mm in front of the rear stop (see attached pictures). I then moved the cage back along the side stop till I hit the ruler.
  • I then clamped the optic down, removed the spacing ruler, and re-adjusted the position of the rear stop to mark the new position of ETMY.
  • We were concerned that the change of position of the cage on the table affected the leveling. Checking with a clean spirit level, we found evidence of a slight tilt in the direction towards the vertex of the IFO, as expected from the way the ETMY cage was moved. To compensate for this, I moved one of the counterweight masses (see attachments) till the spirit level showed the table to be level (to its resolution) in two perpendicular directions
  • We then plugged in the OSEMs into the DB25 connectors on the table. We found that the Y-end electronics were giving different readouts from what we had been seeing in the cleanroom with the X end electronics (not surprising I guess). We resolved to pull out all the OSEMs, check their maximum sensor output values, and re-insert them till the sensor output was half this maximum as best as we could. NOTE TO SELF: UPDATE THE WIKI PAGE!
  • We turned on the damping, and found that the exisiting input matrix performs fairly well.
  • We took a quick look at the spectra of the sensor outputs - interestingly, with the suspension on the seismic stacks inside the chamber, the 16.4 Hz bounce mode peak showed up clearly (these were totally absent in the cleanroom). I did not attempt any fine rotation of the OSEMs in the holders (it is not even clear to me how good/bad the present configuration is) because I reasoned we first need to apply a pitch bias to get the beam back to the ITMY chamber and then re-adjust the OSEM coils. The bounce mode decoupling will be the last step. 
  • For tonight, we decided to leave the optic freely swinging (with EQ stops close by) so that tomorrow, we can look at the offline spectra of sensor outputs and if necessary, re-diagonalize the suspension. 
  • After checking nothing unwanted was left behind in the chamber, we closed it up for tonight.

Plan for tomorrow:

  • Pitch balancing check (by looking at reflected beam at ITMY)
  • Re-adjust OSEMs on ETMY, minimize bounce mode coupling into sensor outputs
  • Make Y arm cavity by re-positioning ITMY

Attachments:

Attachment #1: Wire is in groove in side without OSEM

Attachment #2: Wire is in groove in side with OSEM (picture taken with OSEM coil removed)

Attachment #3: UL magent relative to OSEM coil

Attachment #4: LL magent relative to OSEM coil

Attachment #5: LR magnet relative to OSEM coil

Attachment #6: UR magnet relative to OSEM coil

Attachment #7: Side magnet relative to OSEM coil

Attachment #8: ETMY HR face with F.C. film removed. Non-covered part isn't super clean, but the covered part itself does not have any large specks of dust visible.

Attachment #9: Scheme adopted to shorten Y arm length by 19mm.

Attachment #10: Current situation inside EY chamber. Counterweight that was moved to balance the table is indicated.

 

Attachment 1: IMG_3025.JPG
IMG_3025.JPG
Attachment 2: IMG_3035.JPG
IMG_3035.JPG
Attachment 3: IMG_3030.JPG
IMG_3030.JPG
Attachment 4: IMG_3029.JPG
IMG_3029.JPG
Attachment 5: IMG_3028.JPG
IMG_3028.JPG
Attachment 6: IMG_3027.JPG
IMG_3027.JPG
Attachment 7: IMG_3026.JPG
IMG_3026.JPG
Attachment 8: IMG_3036.JPG
IMG_3036.JPG
Attachment 9: IMG_3038.JPG
IMG_3038.JPG
Attachment 10: IMG_3045.JPG
IMG_3045.JPG
  12433   Tue Aug 23 17:05:20 2016 PrafulUpdateElectronicsSoldered Circuit Working

I remade another soldered circuit, adding extra 100uF electrolytic bypass capacitors at the input and output of the voltage regulator and ensuring that every grounded component now has its own path to ground rather than going through other elements. This circuit now seems to be working just like the solderless circuit. Attached is the transfer function of the soldered circuit, which matches with the result from the solderless circuit.

 

soldered_transfer_function.png

solderless_transfer_function.png

Here are both on the same figure- they are about overlapping but are slightly different if you zoom in enough.

both_transfer.png

I have also attached a new version of the circuit schematic to reflect the changes and to make the physical layout more clear.

simple_ampv2.pdf

My next step for these last few days this summer will be designing a PCB using Altium. I've emailed Varun about how to use Altium on the iMac but he hasn't responded. If anyone else knows how to use the software, please let me know.

Attachment 2: soldered_transfer_function.png
soldered_transfer_function.png
Attachment 3: soldered_transfer_function.png
soldered_transfer_function.png
Attachment 5: solderless_transfer_function.png
solderless_transfer_function.png
Attachment 6: both_transfer.png
both_transfer.png
Attachment 8: both_transfer.png
both_transfer.png
Attachment 10: simple_ampv2.pdf
simple_ampv2.pdf
  12432   Tue Aug 23 09:50:17 2016 Max IsiUpdateGeneralSummary pages down due to cluster maintenance

Summary pages down today due to schedulted LDAS cluster maintenance. The pages will be back automatically once the servers are back (by tomorrow).

  12431   Mon Aug 22 18:35:16 2016 PrafulUpdatePEMthe lab temp is up

The temperature is decreasing slowly but is still above 24 C.

temp_plot.png

Quote:

The IFO room temp is up a bit and it is coming down. The out side temp is not really high.

 

Attachment 1: temp_plot.png
temp_plot.png
Attachment 3: temp_plot.png
temp_plot.png
  12430   Mon Aug 22 18:04:24 2016 gautamUpdateSUSETMY OSEMs inserted

[Johannes, gautam]

We worked on trying to insert the OSEMs in the optimal positions such that the coupling of the bounce mode into the OSEM sensor signals was minimised.

First, I gave the barrel of the optic a wipe with some optical tissue + acetone in order to remove what looked like some thin fibres of dried first contact. It may be that while I was applying the F.C., the HEPA air flow deposited these on the barrel. In any case, they came off easily enough. There is still a few specks of dust on various parts of the barrel, but it is likely that these can just be removed with the ionized air jet, which we can do after putting the optic in the chamber.

We then did the usual OSEM insertion till the magnets neutral position was such that the sensor output was ~50% of the fully open value (turned the HEPA off for the remainder of this work). I tweaked the bottom OSEM plate a little in order to center the magnets relative to the coil as best as possible. Once this was done, we attempted to look at spectra of the sensor outputs, with 0.05 Hz bandwidth - however, we were unable to identify any peak at 16.4 Hz, which is what a Jan 2015 measured value wiki page claims the bounce mode frequency is (although this was an in vacuum measurement). There were a couple of peaks at ~15.7 Hz and ~16.7 Hz, but I can't think of any reason why the bounce mode resonance should have changed so much - after all, this is ETMY for which no standoff regluing was done. The only difference is that there is some first contact + peek mesh on the HR face now, but I doubt this can modify the bounce resonance frequency so much (this is just my guess, I will have to back this up with a calculation).

Anyways we decided to take this up again tomorrow. Things are progressing fairly well now, I hope to be able to put in ETMY back into the chamber at some point tomorrow and commence re-alignment of the interferometer. I've left the OSEMs in for today, with the EQ stops not engaged but close by. HEPA has been turned back on.

  12429   Mon Aug 22 16:33:32 2016 SteveUpdatePEMthe lab temp is up

The IFO room temp is up a bit and it is coming down. The out side temp is not really high.

Attachment 1: temp.png
temp.png
  12428   Mon Aug 22 13:06:11 2016 gautamUpdateSUSETMY suspended

Today morning, I suspended ETMY and made the same checks dscribed below. The clamping went smoothly, 5 in. lb. of torque seems sufficient, in the limited observation time, there has been no evidence of wire sag. Today afternoon, we will go about putting the OSEM coils in, setting their equilibrium points etc. This may need to be re-done once the optic is in the chamber and the first contact has come off, but at least we can coarsely place them in the relative convenience of the cleanroom. 

GV EDIT 9.15pm 22 Aug: Eric had a look at both towers and pointed out that I had neglected to use washers on the wire stops. After consultation with Steve, I decided that it is not worth it to remove the clamp and re-suspend the optic - it is likely that the current suspension process will have caused new grooves in the suspension block, which will have to be removed, and the sanding process did not work so well last time. In any case, the net effect of this will be that the actual torque with which the clamp is tightened will be slightly different from 5 in. lb., but since there is no evidence that the clamp isn't tight enough / is too tight, I think it is okay to push ahead. 

Quote:
  • ETMX has been successfully suspended
  • I've used one of the new wire clamps, and also the new suspension wire 
  • Because the HR face has first contact, pitch balancing cannot be checked at this point. But since the pitch balance was checked after the standoff was glued, there is no reason to believe it would have changed
  • Heights of the two scribe lines were checked with the microscope and verified to be at 5.5" above the tabletop. Also checked the position of the scribe line on the bottom of the optic to make sure the optic wasn't somehow rotated
  • Checked that wire was in the groove in the standoff on both sides, and that the optic was freely hanging with no EQ stops engaged. I also verified that there are no obvious kinks/other funny features where the wire is in contact with the optic barrel below the standoffs.
  • Wire clamps were tightened with the new torque wrench and 5 in. lb. (0.56 N m) of torque. Primary clamp was successfully tightened. However, the wire snapped between the primary and secondary clamps on one side. It is unclear to me how or why this happened. But since the primary wire clamp is the important one, I don't think it is worth re-suspending ETMX all over again
  • I've left the cage on the flow bench for now, with EQ stops engaged. OSEM coils have yet to be inserted, but I suppose we want to do this in the vacuum chamber now to do the fine rotation to minimize the bounce mode in the OSEM signals
  • I've prepared ETMY and its cage for suspension, will work on it tomorrow

 

Attachment 1: IMG_3019.JPG
IMG_3019.JPG
  12427   Sun Aug 21 17:21:22 2016 PrafulUpdateElectronicsProblems with PCB Circuit

For the past week, I've been trying to make a soldered amplifier circuit to use in a prototype box, However, I've been running into this same issue. The circuit, pictured below, works fine on a solderless breadboard.

simple_amp.png

When I amplify a sine wave, I get a clean looking result at the output on the solderless breadboard:

However, on my soldered circuit, if I turn up the negative voltage supply from the power supply past about -12.5V (the target is -15V), I get a strange signal that Gautam suggested looks like some kind of discharging.

At -12.3 V (soldered breadboard):

At -15.0 V (soldered breadboard):

The signal is much noisier. Zooming in on this second signal, this pattern appears:

This pattern is also showing up even when there is no input from the function generator and the circuit is just given a voltage supply of +/- 15V:

I have tried switching out both the positive and negative voltage regulators, the opamp, and remaking and resoldering the entire circuit but I'm still getting the same signal, which is absent from the solderless circuit. This output was produced with a function generator, so I have also ruled out the microphone as a source of this extra noise. The voltage dependence of this problem made me think it was the voltage regulator, but I've switched out the voltage regulator multiple times and it's still showing up. I'm not sure why this signal appears only as the negative voltage supply is increased- there is no problem with increasing the positive input voltage. Please let me know if you have any ideas as to what component or issue could be causing this.

Attachment 2: simple_amp.png
simple_amp.png
Attachment 4: clean.jpg
clean.jpg
Attachment 5: -12.jpg
-12.jpg
Attachment 6: -15.jpg
-15.jpg
Attachment 7: pat1.jpg
pat1.jpg
Attachment 8: pat2.jpg
pat2.jpg
Attachment 10: bad.jpg
bad.jpg
Attachment 11: pattern.jpg
pattern.jpg
Attachment 12: pattern2.jpg
pattern2.jpg
Attachment 13: pat2.jpg
pat2.jpg
Attachment 16: patternzoomed.jpg
patternzoomed.jpg
  12426   Sun Aug 21 16:23:05 2016 gautamUpdateSUSETMX suspended
  • ETMX has been successfully suspended
  • I've used one of the new wire clamps, and also the new suspension wire 
  • Because the HR face has first contact, pitch balancing cannot be checked at this point. But since the pitch balance was checked after the standoff was glued, there is no reason to believe it would have changed
  • Heights of the two scribe lines were checked with the microscope and verified to be at 5.5" above the tabletop. Also checked the position of the scribe line on the bottom of the optic to make sure the optic wasn't somehow rotated
  • Checked that wire was in the groove in the standoff on both sides, and that the optic was freely hanging with no EQ stops engaged. I also verified that there are no obvious kinks/other funny features where the wire is in contact with the optic barrel below the standoffs.
  • Wire clamps were tightened with the new torque wrench and 5 in. lb. (0.56 N m) of torque. Primary clamp was successfully tightened. However, the wire snapped between the primary and secondary clamps on one side. It is unclear to me how or why this happened. But since the primary wire clamp is the important one, I don't think it is worth re-suspending ETMX all over again
  • I've left the cage on the flow bench for now, with EQ stops engaged. OSEM coils have yet to be inserted, but I suppose we want to do this in the vacuum chamber now to do the fine rotation to minimize the bounce mode in the OSEM signals
  • I've prepared ETMY and its cage for suspension, will work on it tomorrow
Attachment 1: IMG_3018.JPG
IMG_3018.JPG
  12424   Fri Aug 19 22:51:12 2016 gautamUpdateSUSETMs first-contacted

I've applied first contact to both the ETMs. They're now ready to be suspended. I've also cut up some lengths of the new wire and put them in the oven for a 12 hour 70C bake. 

  • For both ETMs, I first applied first contact to the bulk of the HR and AR surfaces (all the way out to the edge for the HR, for the AR as large an area as possible without getting too close to the magnets). Calum recommended pouring first contact onto the horizontal optic, but since I had no practise with this method, I opted not to try it out for the first time on our ETMs
  • After allowing this to dry for 24 hours, I peeled this layer off. Visual inspection suggests that the whole film came off cleanly. 
  • I then applied first contact to a smaller area around the center of the optic for only the HR surface. This will only be peeled off once the suspended optic is back in the vacuum chamber. This way, we keep the HR face protected for as long as possible.
  • Even though we applied F.C to both faces of the ITMs, I don't think its so important to keep a film on the AR side of the ETMs till we take it in. So I didnt re-coat the AR side with a smaller area of F.C. This way, if we want, we can do the OSEM assembly in the cleanroom without having to worry about peeling the F.C off with limited access to the rear of the optic.
  • I also opted to bake some lengths of the newly arrived steel wire for suspension. Not sure how important/useful this bake will be.

Unless we want the AR surface to also have a small F.C coat until the optic is in the vacuum chamber, I think I will proceed with re-suspending the ETMs..

  12423   Thu Aug 18 15:16:09 2016 SteveUpdateSUSSOS sus wire is in

 

Stress Relieved 0.0017"  Music Wire  CFW P/N: CFW2035025,   Made 08-17-2016                                

    Old  2003    New  2016  
GBL 358.9 240.610 grams
UTS 357,061 229,603 PSI
YTS 343,211 177,371 PSI
ELONG 2.38 0.8 %
HEAT 10622 10622  

GBL (grams breaking load )         

UTS (ultimate tensile strength)                                 

YTS (yield tensile strength) 

ELONG (elongation)  

Quote:

0.0017" OD., 500ft steel music wire ordered. Pictures of the existing roll are below. It will be on 8" OD. spool too.

 

 

Attachment 1: 0.0017new.jpg
0.0017new.jpg
  12422   Thu Aug 18 14:14:20 2016 gautamUpdateSUSAir-bake of towers - finished

I took the two cages, wires and wire clamps out this morning, back into the cleanroom after their 12 hour 70C bake. 

I've also applied first contact to the AR face of the optics. Steve is preparing a jig which will allow us to apply first contact on the HR side with the optic horizontal. The idea is to apply a large coating first, to clean the bulk of the HR surface, and peel it off before re-suspending the optic. Then we can paint on a smaller area, suspend the optic (and hope the pitch balancing is alright) before taking the whole assembly into the chamber where it will be peeled off. 

Calum recommended that we buy a new ionizing gun + electrometer assembly (apparently our current set up is woefully obsolete) but I don't know if we can have these in time for the first contact peeling...

  12421   Thu Aug 18 08:17:16 2016 SteveUpdateSUSwire clamping preparation

The wire inprints were removed by 800P grain paper [Norton 73568] The SS bridge block now has an undesireble vally in the wire location.

The sus bridges were soaked in acetone over night and sonicated to remove residual sand paper.

 

Quote:

I just put in the following into the air bake oven for a 12 hour, 70C bake:

  • ETMX and ETMY cages (with sanded suspension blocks loosely tightened for now, we will tighten them after the bake)
  • 13 new wire clamps that were recently made by the shop
  • 7 lengths of suspension wire (since the new wire is unlikely to arrive for another 2 weeks). This should be sufficient in case we overtighten the wire clamps a couple of times and the wire snaps.

I put these in at 10.30pm. So the oven will be turned off at 10.30am tomorrow morning. The oven temperature seems stable in the region 70-80 C (there is no temperature control except for the in built oven control, I just adjusted the dial till I found the oven remains at ~70C.

Tomorrow, we will look to put on first contact onto the ETMs, and then get about to re-suspending them.

 

Attachment 1: 800P.jpg
800P.jpg
  12420   Wed Aug 17 23:00:57 2016 gautamUpdateSUSAir-bake of towers

I just put in the following into the air bake oven for a 12 hour, 70C bake:

  • ETMX and ETMY cages (with sanded suspension blocks loosely tightened for now, we will tighten them after the bake)
  • 13 new wire clamps that were recently made by the shop
  • 7 lengths of suspension wire (since the new wire is unlikely to arrive for another 2 weeks). This should be sufficient in case we overtighten the wire clamps a couple of times and the wire snaps.

I put these in at 10.30pm. So the oven will be turned off at 10.30am tomorrow morning. The oven temperature seems stable in the region 70-80 C (there is no temperature control except for the in built oven control, I just adjusted the dial till I found the oven remains at ~70C.

Tomorrow, we will look to put on first contact onto the ETMs, and then get about to re-suspending them.

Attachment 1: IMG_3006.JPG
IMG_3006.JPG
  12419   Wed Aug 17 22:09:04 2016 ranaUpdateSUSSOS sus wire ordered

Not really true that it passed. That's just an arbitrary margin. Best to throw away all the old wire. We have no quantitative estimate of what the real torque should be. Its just feelings.

Quote:

The wire will arrive in 1-2 weeks. It is a new production. Brad Snook of Ca Fine Wire was suprised that we are still using the 13 years old wire.  Oxidation is an issue with iron contained steel wire.

He would not give me a shelf life time on it. He recommended to check the strenght of it before usage. It passed with safety factor of 2 just recently.

In the future we'll store the new spool in oxigen free nitrogen environment..

 

  12418   Wed Aug 17 16:28:46 2016 KojiUpdateCOCRC folding mirrors - Numerical review

For the given range of the PR3/SR3 RoCs for both cases, all the resulting numbers such as TMSs/mode matching ratios look reasonable to me.

  12417   Wed Aug 17 14:37:36 2016 gautamUpdateCOCRC folding mirrors - Numerical review
Quote:

 

Cavity One-way Gouy phase [rad]           TMS [MHz]           
PRX 0.244 1.730
PRY 0.243 1.716
SRX 0.197 1.743
SRY 0.194 1.717

So, there are regions in parameter space for both options (i.e. keep current G&H mirrors, or order two new sets of folding mirrors) that get us close to the design numbers...

Keeping these design numbers in mind, here are a few possible scenarios. The "designed" TMS numbers from my previous elog are above for quick reference.

Case 1: Keep existing G&H mirror, flip it back the right way, and order new PR3/SR3. 

  • Spec PR3 to be concave with RoC 600 +/- 50m
  • This means the TMS in the PRC is in the range 1.4 MHz - 1.6 MHz [see this plot]
  • The mode matching efficiency for the PRC is > 98.5% [see this plot]
  • The TMS in the SRC is in the range 1.6 MHz - 1.8 MHz [see this plot]
  • Mode matching efficiency for SRC is > 98.5% [see this plot]
  • PRG between 34-38, depending on uncertainty in measurement of RoC of existing G&H mirror [see Attachment #1, added Nov 11 2016]

Case 2: Order two new sets of folding mirrors

  • Spec PR3/SR3 to be flat - for purposes of simulation, let's make it concave with RoC 10 +/- 5 km
  • Spec PR2/SR2 to be concave with RoC 1500 +/- 500m
  • The TMS in the PRC is between 1.7 MHz and 1.85 MHz [see this plot]
  • Mode matching efficiency is >98.5% in the PRC [see this plot]
  • TMS in the SRC is between 1.7 MHz and 2 MHz [see this plot]
  • Mode matching efficiency >99.0% in the SRC [see this plot]

At first glance, it looks like the tolerances are much larger for Case 2, but we also have to keep in mind that for such large RoCs in the km range, it may be impractical to specify as tight tolerances as in the 100s of metres range. So these are a set of numbers to keep in mind, that we can re-iterate once we hear back from vendors as to what they can do.

For consolidation purposes, here are the aLIGO requirements for the coatings on the RC folding mirrors: PR2, PR3, SR2, SR3

Attachment 1: PRG.pdf
PRG.pdf
  12416   Wed Aug 17 08:47:43 2016 ericqUpdateSUSAir-bake finished

I turned off the air bake oven at 8:45AM. I'll leave the optics alone for a bit while it cools.

  12415   Tue Aug 16 21:54:27 2016 gautamUpdateSUSAir-bake - IN PROGRESS

I put in both ETMX and ETMY into the air-bake oven at approximately 8.45pm tonight. They can be removed at 8.45am tomorrow morning. 


  • Given that we had previously melted a thermocouple in this oven, and there have been no high temperature bakes in it since, we ran the oven at 100C for about 3 hours in the afternoon
  • After that, I left the oven door open for an hour for the interior to return to room temperature
  • I then re-connected the controller (which doesn't seem very precise, it pulses the AC power to the oven in order to control the temperature), and dialled the oven back down to heating level 4, which is what Bob had it set at. I then waited for a couple of hours for the oven to reach ~34C
  • Before putting the optics in, I gave the inside of the oven a quick wipe with a clean wipe, and palced a layer of Al foil on the bottom of the oven
  • The optics are sitting on their donuts (see Attachment #1) - the copper wire elevates the optic+donut slightly and provides a path for air flow
  • ETMY was drag wiped with acetone+isopropanol prior to baking (to remove acetone stains from soaking to remove epoxy residue
  • We will of course be cleaning the optics with first contact prior to re-installation in the vacuum chambers
  • I am not sure what the extra cylindrical piece in there is, but Bob advised me to leave it in there so that's what I did
  • I've observed the temperature over ~2hours since I first put it in, and the oven/controller isn't going bonkers, so I'm trusting the controller and leaving for the night
Attachment 1: IMG_3005.JPG
IMG_3005.JPG
  12414   Tue Aug 16 16:38:00 2016 gautamUpdateCOCRC folding mirrors - Numerical review

Here are the results for case 2: (flat PR3/SR3, for purpose of simulation, I've used a concave mirror with RoC in the range 5-15km, and concave PR2/SR2 - I've looked at the RoC range 300m-4km).

  • This is where we order two new sets of mirrors, one for use as PR2/SR2, and the other for use as PR3/SR3.
  • RoC of flat PR3/SR3 in simulation explored in the range 5km-15km (concave)
  • RoC of concave PR2/SR2 in simulation explored in the range 300m-4km (concave)

Attachment #1: Mode matching between PRC cavities and arm cavities with some contour plots

Attachment #2: Mode matching between SRC cavities and arm cavities with some contour plots

Attachment #3: Gouy phase and TMS for the PRC. I've plotted two sets of curves, one for a PR3 with RoC 5km, and the other for a PR3 with RoC 15km

Attachment #4: Gouy phase and TMS for the SRC. Two sets of curves plotted, as above.


Hopefully EricG will have some information with regards to what is practical to spec at tomorrow's meeting.


EDIT: Added 9pm, 16 Aug 2016

A useful number to have is the designed one-way Gouy phase and TMS for the various cavities. To calculate these, I assume flat folding mirrors, and that the PRM has an RoC of 115.5m, SRM has an RoC of 148m (numbers taken from the wiki). The results may be summarized as:

Cavity One-way Gouy phase [rad]           TMS [MHz]           
PRX 0.244 1.730
PRY 0.243 1.716
SRX 0.197 1.743
SRY 0.194 1.717

So, there are regions in parameter space for both options (i.e. keep current G&H mirrors, or order two new sets of folding mirrors) that get us close to the design numbers...

Attachment 1: PRC_consolidated.pdf
PRC_consolidated.pdf
Attachment 2: SRC_consolidated.pdf
SRC_consolidated.pdf
Attachment 3: GouyPRC.pdf
GouyPRC.pdf
Attachment 4: GouySRC.pdf
GouySRC.pdf
  12413   Tue Aug 16 11:51:43 2016 gautamUpdateCOCRC folding mirrors - Numerical review

Summary of roundtable meeting yesterday between EricG, EricQ, Koji and Gautam:

We identified two possible courses of action.

  1. Flip the G&H mirror (PR2/SR2) back such that the (convex) HR face is the right way round. We want to investigate what are the requirements on a new PR3/SR3 optic that will guarantee cavity stability and also give good mode matching.
  2. Order two new sets of mirrors (i.e. replace all 4 folding mirrors). In this case, we want to spec a flat (how flat is reasonable to specify? EricG will update us) PR3/SR3, and design a PR2/SR2 with some concavity that will guarantee cavity stability in the event PR3/SR3 deviates from flatness (but still within what we spec). The choice to make PR3 as close to flat as possible is because the angle of incidence in our arrangement means that any curvature on PR3 dominates astigmatism.

I have done some calculations to evaluate the first alternative. 

  • Based on yesterday's preliminary discussion, we felt it is not reasonable to spec mirrors with RoC > 4km (sag of ~80nm). So I restrict my analyses to the range 300m-4km
  • Koji has a measurement of the phase maps for the G&H mirrors. The measured curvature is ~-500m. In my simulations, I've tried to allow for error in this measurement, so I look at the range -450m to -700m for the G&H mirror.
  • The Gouy phase analysis suggests we should look for an RoC of +500m (concave) for the new PR3/SR3 to have a TMS of ~1.5 MHz. Anything flatter (but still concave) means the TMS gets smaller.
  • The mode-matching in this region also looks pretty good, between 98% and 99%
  • I will post results of the analysis for the second alternative here for comparison

Something else that came up in yesterdays meeting was if we should go in for 1" optics rather than 2", seeing as the beam spot is only ~3mm on these. It is not clear what (if any) advantages this will offer us (indeed, for the same RoC, the sag is smaller for a 1" optic than a 2").


Attachments:

Attachment #1: Mode-matching maps between PRX and Xarm cavities, PRY and Yarm cavities with some contours overlaid.  

Attachment #2: Mode-matching maps between SRX and Xarm cavities, SRY and Yarm cavities with some contours overlaid. 

Attachment #3: Gouy phase calculations for the PRC

Attachment #3: Gouy phase calculations for the SRC

 

Attachment 1: PRC_consolidated.pdf
PRC_consolidated.pdf
Attachment 2: SRC_consolidated.pdf
SRC_consolidated.pdf
Attachment 3: GouyPRC.pdf
GouyPRC.pdf
Attachment 4: GouySRC.pdf
GouySRC.pdf
  12412   Tue Aug 16 08:04:49 2016 SteveUpdateSUSETMY had Aluminum bridge

The SOS ETMY tower had and Aluminum bridge. How is it possible that this was true? Is SS better than Al for some quantitative scientifc reason ???

Their weight ratio as measured  =  922 / 307 g = 3

Destinies:  SS 304 / Al 6061  >  0.289 / 0.098  [ lb /in3 ] =  2.94

Attachment 1: AlBridge.jpg
AlBridge.jpg
  12411   Mon Aug 15 18:28:15 2016 gautamUpdateSUSAir-bake preparation

I assume that we are prepared to live with the pitch bias situation of ETMY (i.e. we can achieve a configuration in which there is some pitch bias to the coils, and the OSEMs are inserted such that the PD outputs are half their maximum value). Or at least that we don't want to go through the whole standoff-regluing procedure for ETMY as well.

So today I took the optic out, and began to make some preparations for the air bake.

  • Both optics are now sitting in their respective metal donuts. 
  • How do we want to bake the optics? Bob has said he has prepared the oven for this bake, and that he has configured the temperature controller to a setpoint of 34C, and a ramp time of 2 hours to reach that temperature from lab temperature (we should check this before putting the optics in there with our independent temperature sensor - also, he is away for the week now so we can't get his input on any of these). But what about the actual logistics of how the optics are going to be housed? Specifically:
    • Do we want the donut to sit on some sort of tray? Presumably it is not ideal to have the HR surface in close proximity to the oven floor? 
    • Does the oven need any special cleaning?
    • Do we cover the donut+optic setup with a glass jar? If we do, any particles we eject off the optic can't escape the confines of the bowl, and if we don't, detritus from elsewhere may settle on the optic?
    • How long do we want this bake to last? 24hours? 48 hours? Bob didn't have an answer when I asked him earlier in the afternoon...
  • I also removed the suspension block from the top of the towers of both ETMX and ETMY, so that Steve could work on sanding them before we acetone-wipe and bake the towers themselves.
    • It was very apparent that the weights of the two pieces were largely different (ETMY suspension block ~350g, ETMX suspension block ~960g), even though they have the same physical dimensions.
    • Investigation into why this was yielded nothing conclusive. But Steve and I think that the ETMY suspension block is made out of Aluminum rather than SS, which would explain why the wire grooves seem deeper in the ETMY piece than the ETMX piece. It is worth noting that the specification calls for SS and not aluminum. But the top piece of the ETMY suspension (and indeed the old ETMX suspension) looks different from the specification, in that they don't have tapped holes for the secondary wire clamps (see Attachment #1).
    • I'm not sure if this is important, but it is worth noting. Steve and I also checked the remaining suspension towers. We think that ITMY, BS, SRM and PRM have the correct (to specification) suspension block. We couldn't get a look at ITMX and didn't want to take the door off. So ETMY (and possibly ITMX) will be the only suspension(s?) with a different suspension block.
  • Steve's sanding efforts did not go ideally.
    • He was successful in removing the wire grooves.
    • But the sharp edge which is supposed to clamp the wire seems to have been rounded a little bit (see Attachment #1). 
    • Overall, the section that we was sanded looks lower (i.e. its like we've dug a small channel into the plane of the suspension block)
    • Given that we suspect the ETMY suspension block is Aluminum, it is likely that attempting to sand it will yield an even deeper channel.
  • Do we want to bake the suspension towers in the large baking oven? Presumably we don't want to bake the optics with anything else. But does the large oven need any special cleaning before we stick the towers in there?
  • ETMY has some acetone marks on it. I will try and have this removed by drag wiping with more acetone and isopropanol prior to the bake tomorrow. Anyways we will first-contact clean the HR (and AR) sides after the bake before installing the optic.

In summary, the questions that remain (to me) are:

  1. Are we okay using an Al suspension block?
  2. How perfectly do we want wire grooves from prior suspensions removed? It looks like sanding doesn't work well, do we want to consider sending this into the shop?
  3. Baking logistics, as described above.

I think we can start the baking of the optics tomorrow. The timeline for the suspension towers is unclear, depends on how we want to deal with the sanding dilemma.

Attachment 1: IMG_6816.JPG
IMG_6816.JPG
  12410   Mon Aug 15 14:34:33 2016 ericqUpdateSUSSOS sus wire ordered

We have indeed seen numerous tarnished/rusty points along the wires, and just tried to choose lengths free of any of these. I wonder if this can explain the brittleness/ease with which we've been breaking it. My feeling is that we should use the newer wire if feasible.

  12409   Mon Aug 15 14:29:32 2016 SteveUpdateSUSSOS sus wire ordered

The wire will arrive in 1-2 weeks. It is a new production. Brad Snook of Ca Fine Wire was suprised that we are still using the 13 years old wire.  Oxidation is an issue with iron contained steel wire.

He would not give me a shelf life time on it. He recommended to check the strenght of it before usage. It passed with safety factor of 2 just recently.

In the future we'll store the new spool in oxigen free nitrogen environment..

Quote:

0.0017" OD., 500ft steel music wire ordered. Pictures of the existing roll are below. It will on 9" OD. spool too.

 

 

ELOG V3.1.3-